Circular No. 9061

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V5583 SAGITTARII = NOVA SAGITTARII 2009 No. 3 K. Nishiyama, Kurume, Fukuoka-ken, Japan; and F. Kabashima, Mi-

yaki-cho, Saga-ken, Japan, reported their discovery of a possible nova (mag 7.7) on two 60-s unfiltered CCD frames (limiting magnitude 13.3) taken on Aug. 6.494 and 6.495 UT using a 105-mm f/4 lens; their confirming frames taken around Aug. 6.494 with a 0.40-m reflector yield a position α = 18^{h}07^{m}07^{s}.67, δ = −33^{o}46′33′′.9 (equinox 2000.0). Their survey frames on July 22.531 and 29.584 showed nothing at this position (limiting magnitude 12.7), and nothing is visible on Digitized Sky Survey images. Additional CCD magnitudes for the nova: Aug. 6.182, V = 7.78 (G. Pojmanski, D. Szczygiel, and B. Pilecki, Warsaw University Astronomical Observatory, ASAS3V instrument of the All Sky Automated Survey; independent discovery; position end figures 08^{s}, 46′.6); 7.57, V = 7.43, B–V = +0.60, V –R_{c} = +0.48, V –I_{c} = +0.99 (H. Maehara, Kwasan Observatory, Kyoto University; and Y. Sakane and H. Yamaoka, Kyushu University; 25-cm telescope; position end figures 07^{s}.71, 34′′.1); 7.579, 8.3 (L. Elenin, Moscow, Russia, remotely using a Takahashi Epsilon 180 telescope located at Pingelly, Australia; position end figures 07^{s}.67, 34′′.9); 8.122, I_{c} = 6.52 (S. Kiyota, Ibaraki, Japan, 30-cm GRAS telescope near Mayhill, NM, U.S.A.); 8.130, V = 7.68 (Kiyota); 8.131, R_{c} = 7.41 (Kiyota); 8.133, B = 8.07 (Kiyota). Spectroscopy by M. Fujii (Fujii Bisei Observatory, Okayama, Japan; communicated Maehara) on Aug. 7.60 shows the Balmer series (Hα, Hβ, and Hγ) and Fe II emission lines; the FWHM of the H-alpha emission is 2300 km/s. Spectroscopy by A. Arai, T. Komatsu, M. Yamanaka, M. Sasada, and R. Itoh (Higashi-Hiroshima Observatory, Hiroshima University) with the 1.5-m KANATA telescope on Aug. 7.63 shows the Balmer lines (Hα and Hβ) and Fe II emission lines, with Hα showing an asymmetric profile (its FWHM being ∼ 2100 km/s). Additional details were published on CBETs 1899 and 1900.

E. Kazarovets, Moscow, reports that the GCVS team has assigned the designation V5583 Sgr to this nova. Additional visual magnitude estimates, provided in part by E. O. Waagen, AAVSO: Aug. 8.367, 7.8 (Y. Kok, Stanmore, NSW, Australia); 8.388, 7.6 (H. Matsuyama, Kanimbla, Queensland, Australia); 8.497, 7.9 (Maysuyama); 8.553, 8.2 (Kok); 9.06, 8.2 (J. G. de S. Aguiar, Campinas, Brazil); 9.93, 8.5 (Aguiar); 10.93, 8.8 (Aguiar); 11.92, 8.9 (Aguiar); 12.92, 9.1 (Aguiar); 13.92, 9.2 (Aguiar); 14.083, 9.4 (W. Souza, Sao Paulo, Brazil); 14.93, 9.5 (Aguiar); 15.083, 9.9 (Souza); 15.99, 9.8 (Aguiar); 16.104, 9.6 (Souza); 17.125, 9.9 (Souza).

2009 August 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green